10. Two men chop one ton of wood in two hours. How long will it take three men to chop three tons of wood?
(A) 2 hours (B) 5/2 hours (C) 3 hours (D) 10/3 hours (E) None of these

Answer by Theo(13342)

let R = number of tons of wood per hour that one man can chop.
-----
let W = number of tons of wood that are chopped.
-----
let H = the number of hours that it takes to chop W tons of wood.
-----
let M = number of men
-----
general formula would be R*M*H = W
-----
The first statement says that 2 men chop one ton of wood in 2 hours.
-----
M = 2
W = 1
H = 2
-----
formula of R*M*H = W becomes R*2*2 = 1
solving for R gets us R = 1/4
-----
R = 1/4 means 1 man can chop 1/4 ton of wood in one hour.
-----
The second statement asks for how long it takes 3 men to chop 3 tons of wood.
-----
M = 3
W = 3
R = 1/4
-----
formula of R*M*H = W becomes (1/4)*3*H = 3
solving for H gets us H = 4
-----
Answer is 4 hours.
-----
you have 3 men and each of them can chop 1/4 tons of wood per hour.
At the end of 1 hour they have chopped 3/4 tons of wood.
In 4 hours they have chopped 4 * (3/4) tons of wood which equals 3 tons of wood.
-----